This morning at approximately 10:33am our datacentre experienced a fault with the power switching gear due to external power fluctuations. The UPS' protected the site and discharged. Power was lost to all cabinets at this time. 

The NOC are speaking with our vendors and our electrical contractors to identify a cause and prevent this kind of incident in the future.

NOC  technicians were alerted and responded immediately but power had been lost by the time they arrived on site. They immediately worked to restore power and core services. Power was restored at 10:53am. and servers were brought back online shortly later

We apologise for any inconvenience caused, and should you require any assistance please contact us via the helpdesk.


Sunday, November 18, 2012





« Back